OpenStudy (anonymous):

find the maximum speed at which a truck can safely travel without troppling over,on a curve of radius 250m. the height of the centre of gravity of the truck above the ground is 1.5m and distance between the wheel is 1.5m ,the truck being horizantal.

OpenStudy (anonymous):

@Herp_Derp @henpen

OpenStudy (anonymous):

\[Vmax = \sqrt{r g}\]

OpenStudy (anonymous):

Do u have answer key ? @Ruchi.

OpenStudy (anonymous):

no why? @Yahoo!

OpenStudy (anonymous):

i just need to check my answer

OpenStudy (anonymous):

in book answer is given 35 m

OpenStudy (anonymous):

no its answer is 35m/s

OpenStudy (anonymous):

Can u check ur question once more

OpenStudy (anonymous):

i hav check my question is writing u r appling the concept of troopling bt its in ques given the case of not troopling.

OpenStudy (anonymous):

@timo86m @TuringTest plz hlp me out.

OpenStudy (anonymous):

@tryder @Kryten @ujjwal hlp me out plzz

OpenStudy (ujjwal):

AC=1.5 m AB=(1.5/2)=0.75 m angle of inclination, \(\theta\)=90-\(\tan^{-1}\frac{1.5}{0.75}\)=26.56 degree Now, max velocity, v=\(\sqrt{rg\tan\theta}\) where r is normal reaction

OpenStudy (ujjwal):

Sorry, r is radius! And it will give you an answer of 34.996 (approx 35 m/s)

OpenStudy (anonymous):

tnx for hlp me i hav get the answer bt nt by this 1 by applying torque concept.